In an isosceles triangle the base is represented by x and each leg is four times as long as the base. Write and solve an inequality to represent: The perimeter of the isosceles triangle is at least 45

Answer by checkley79(3341)

A+B+C=PERIMETER.
4X+4X+X>=45
9X>45
X=45/9
X=5 ANS. FOR THE BASE.
PROOF:
4*5+4*5+5=45
20+20+5=45
45=45